# Cosmological Zoom-In Simulation of Odd Radio Circles as Merger-Driven Shocks in Galaxy Groups

### v1

@@ Abstract
Context. A new class of distinct radio objects has been recently discovered, which are now commonly referred to as Odd Radio Circles (ORCs). The origin of these features is yet unclear because their peculiar properties are a challenge for our current understanding of astrophysical sources for diffuse radio emission.
@@
Methods. By modeling the assembly of a massive galaxy group with a final virial mass of in a magnetohydrodynamic zoom-in simulation with on-the-fly cosmic ray treatment, we are able to derive the X-ray and radio properties of the system self-consistently and compare them to observations.
Results. We show that the X-ray properties for the simulated system are agreeing with characteristics of observed galaxy groups in the regarded mass range, legitimating the comparison between the radio properties of the simulated halo and those of observed ORCs. A major merger between two galaxies in the simulation is triggering a series of strong shocks in the circumgalactic medium, which in unison are forming a ring if the line of sight is perpendicular to the merger axis. The shock is rapidly expanding in radial direction and quickly reaches the virial radius of the halo. This formation channel can hence readily explain the morphology and large extent of ORCs. However, the inferred radio luminosity of these features is lower than for observed counterparts, while the degree of polarization seems to be systematically overpredicted by the simulation.
Conclusions. Fossil cosmic ray populations from AGN and stellar feedback might be necessary to explain the full extent of the radio properties of ORCs, since diffusive shock acceleration was the only source term for non-thermal electrons considered in this work.
@@
The current advent of next-generation radio telescopes is delivering a number of interesting surprises to the astrophysics community. Their unprecedented depth and frequency coverage continues to reveal stunning details in various extraterrestrial settings – from galactic sources such as supernova remnants (e.g. Weiler et al. 1986; Dokara et al. 2021; Ball et al. 2023; Do et al. 2024) over to the plentiful extragalactic zoo of objects like star forming disc galaxies (e.g. Martinsson et al. 2016; Krause et al. 2018; Stein et al. 2023; Heesen et al. 2024), AGN-hosting galaxies (e.g. McKean et al. 2021; Venturi et al. 2022; Koribalski et al. 2024a), as well as halos and radio relics of galaxy clusters (e.g. van Weeren et al. 2010; Rajpurohit et al. 2020; Brüggen et al. 2021; Venturi et al. 2022; Cassano et al. 2023; Koribalski et al. 2024c). However, it seems that yet another kind of distinct source needs to be added to the list of possible radio objects. As their name suggests, Odd Radio Circles (ORCs) are peculiar radio features with edge-brightened ring- or shell-like morphology (Norris et al. 2021; Koribalski et al. 2021, 2024b). The spectral indices measured in these features, that can lie between and (assuming that the emission spectrum follows ), suggest that these regions are filled with shock-accelerated cosmic rays (CRs), emitting via synchrotron in radio. When found at low galactic latitudes (e.g., Koribalski et al. 2024c) it can be difficult to identify a host galaxy, leaving the possibility that these radio rings are supernova remnants in or near our Galaxy (Filipović et al. 2022; Omar 2022a; Sarbadhicary et al. 2023). Bonafide ORCs, however, have a central early-type galaxy, strongly suggesting they are of extragalactic origin (Koribalski et al. 2021; Norris et al. 2022, 2025). Since ORCs lack a counterpart in any other wavelength regime, distance and size measurements of ORCs rely on the redshifts of their putative host galaxies. Judging by the elliptical galaxies visible in optical that are situated near the geometrical centers of ORCs, these objects lie at distances of and often have diameters of several hundred kpc, while the stellar masses of the central galaxies lie around . This means that ORCs have enormous sizes compared to their central host, as their extent can even exceed the virial radius of the host system.
@@
Since the nature of ORCs is still under debate, one needs to be careful in controlling the sample of considered objects, in order to ensure an appropriate comparison. The following set of conditions seems to works best as a common denominator: the ORC has to a) display a ring or shell-like morphology with b) at least one elliptical galaxy in its center, where c) the size of the radio ring is large, i.e. comparable to the virial radius of the central galaxy (group). While the neighbouring ORCs 2 & 3 from the original discovery paper by Norris et al. (2021) – who first reported such radio morphologies and have introduced the term ORC – are likely the lobes of a re-started radio galaxy where one lobe has ring morphology, there is a growing list of different ORCs whose origin is yet unclear. It is important to note, that the observed morphologies can depend on the resolution and frequency of the observation facility. While an object may look diffuse in one image, the characteristic ring emission for ORCs can reveal itself in surveys with higher resolution and sensitivity. The following observed objects are falling under the criteria stated above: ORCs 1 & 4 (Norris et al. 2021), ORC 5 (Koribalski et al. 2021), ORC 6 (Dolag et al. 2023; Koribalski et al. 2025), ORC J2223–4834 (Gupta et al. 2022), SAURON (Lochner et al. 2023), Physalis (Koribalski et al. 2024b), Cloverleaf (Bulbul et al. 2024), ORC J1027–4422 (Koribalski et al. 2024c) and ORC J0219–0505 (Norris et al. 2025). There are several ongoing searches for ORCs and similar systems as for example the study by Gupta et al. (2025).
@@
A variety of extragalactic formation scenarios for ORCs have been proposed. Several works have addressed the scenario of a termination shock due to galactic outflows during starburst events (Norris et al. 2021, 2022; Coil et al. 2024). Then there is the possibility that a black hole is responsible for these features, either via binary mergers of supermassive black holes (Koribalski et al. 2021; Norris et al. 2022), tidal disruption of stars by black holes (Omar 2022b), as well as AGN jets, where the ejected radio lobes can adopt a circular appearance under certain projections and physical conditions (Nolting et al. 2023; Fujita et al. 2024; Lin & Yang 2024). In this regard, Shabala et al. (2024) have pointed out that the radio power of a mere AGN jet would decline too quickly to explain the properties of ORCs, and hence concluded that an additional shock running through preexisting AGN bubbles would be necessary in this scenario. Finally, there is the shock-driven formation channel that was first proposed by Dolag et al. (2023) and Koribalski et al. (2024b). Based on cosmological simulations, the suggested concept was that merger-accelerated shocks – i.e. merger shocks of coalescence events in massive galaxies and groups that run into outer accretion shocks – can produce features that very closely resemble ORCs. Such processes have been studied for a while now in relation to galaxy clusters, where merger shocks are most likely source of for radio relics (e.g. Vazza et al. 2009; Ha et al. 2018; Zhang et al. 2019). But in the case of clusters, these features always lie well within the boundaries of the respective halo. Hence, ORCs could be unprecedented evidence for merger shocks in galaxy and group halos, where their extent is pushing beyond the boundaries of the virial radius. However, the simulations by Dolag et al. (2023) predicted radio luminosities that were too low compared to observations, concluding that the relatively low simulated halo mass of could be a major issue, since the total energy budget is underestimated. This problem was also pointed out by Yamasaki et al. (2024), who explored the feasibility of pure accretion shocks as the source for ORCs analytically, concluding that more complex simulations with a larger halo mass of at least are required, since these are the estimated halo masses of the ORC host galaxies.
@@
In the present work, we are addressing the necessity of such a study. We present a first of its kind zoom-in simulation of a galaxy merger in high resolution, where the final group mass is corresponding to observed halos in ORCs. We performed a non-radiative magneto-hydrodynamic simulation with an on-the-fly treatment of spectral CR electrons. First, we summarize the numerical details of the simulation by stating the important features of the code and initial condition in Sect. 2. In Sect. 3, we give a detailed presentation of the simulation results. We outline the time evolution of the merger event and then describe the halo’s X-ray and radio properties, providing inferred mock observation images. Hereby, a particular focus lies on the polarization of the simulated ORC. Similarities and differences with observations are discussed in Sect. 4, before we summarize our results in Sect. 5.
@@ 2 Simulation code and setup
The simulation was carried out with the cosmological code OpenGadget3, which is utilizing a smoothed particle magnetohydrodynamics (SPMHD) scheme to model gas dynamics and a Tree-PM solver for the gravity (release paper in prep.). Its infrastructure is based on its predecessor Gadget2 (Springel 2005) but employs major improvements in the gravity solver (Ragagnin et al. 2016) and SPH implementation (Beck et al. 2016b), where also a revised time step criterion is ensuring the stability of the code around strong shocks in a simulation (Pakmor et al. 2012). We use the Wendland C4 function as the kernel to reconstruct the local hydrodynamic properties of the simulated fluid based on the position and smoothing length of the simulated particles (Dehnen & Aly 2012; Donnert et al. 2013). The code includes both a physically motivated conductivity (Jubelgas et al. 2004; Dolag et al. 2004; Petkova & Springel 2009), as well as an artificial conduction and time-dependent viscosity scheme (Price 2012; Beck et al. 2016b), which provides numerical convergence when compared to other hydrodynamic discretization methods.
@@
The code resolves the magnetohydrodynamic equations in Lagrangian form (Dolag et al. 1999; Dolag & Stasyszyn 2009) in the non-ideal regime, i.e. with non-vanishing resistivity (Bonafede et al. 2011). Non-zero contributions of the magnetic field , that can emerge due to numerical truncation errors, are suppressed by the hyperbolic divergence cleaning method described by Tricco et al. (2016), where the implementation into OpenGadget3 will be described by Steinwandel & Price (in prep.). We run the code with an on-the-fly shock finder (Beck et al. 2016a), that is returning the shock’s hydrodynamic quantities, as well as as the Alfvénic Mach number and shock obliquity (Böss et al. 2023b). Finally, we include cosmic ray electrons with the module CRESCENDO, which is a Fokker-Planck solver for spectral cosmic rays (CRs, Böss et al. 2023b). The subgrid model computes and evolves the distribution function of CRs for each resolution element (i.e. SPH particle) at each time step in the simulation. The source for CRs is diffusive shock acceleration (DSA) – which therefore heavily relies on the output of the built-in shock finder – and applies shock obliquity dependent acceleration efficiencies (Pais et al. 2018; Kang 2024). These injected populations can then further evolve in time via adiabatic losses as well as radiative losses, which are inverse Compton scattering and synchrotron emission for CR electrons (for the computation of the synchrotron emissivity, c.f. Appendix 2 by Böss et al. 2023a). The present simulation was performed in a non-radiative regime, which does not include star formation, but only resolves the dark matter and gas evolution.
@@
The initial condition for the simulation is drawn from the COMPASS111http://www.magneticum.org/complements.html#Compass set. It is a suite of zoom-in setups for a diverse selection of halos, from massive clusters down to Milky Way-like galaxies, which were selected from a dark matter-only cosmological simulation with a boxlength of (e.g. Bonafede et al. 2011). We select the dfrogin region, which produces a galaxy group halo with a final virial mass of (Schlachtberger 2014). The gas mass resolution and smoothing length of the simulation are listed in Table 1. The simulation employs the standard WMAP-7 CDM cosmological parameters (Komatsu et al. 2011).
@@ 3 Results
Until , the halo undergoes an assembly process that includes various mergers with different galaxies. Shortly after , the main halo experiences a particularly violent event, with a mass ratio of . It is this major merger that is triggering the shock and CR injection with subsequent radio emission that we identify as an ORC in this paper. In the following sections we will analyze the time evolution and properties of the galaxy group around this merger event.
@@ 3.1 Gas and shock morphology
Fig. 1 shows the evolution of the system around the time of the major merger event. The rows represent three different output times which are separated by . The left and right halves of the figure display the X-ray surface brightness in the band and the cosmic ray electron energy density, respectively. The X-ray flux was estimated based on contributions from thermal and line emission (assuming ), using the software Smac (Dolag et al. 2005a). We assume a distance of (corresponding to ), as it is the case for the observed ORC Physalis (Koribalski et al. 2024b). We show two projections in each case, which show the system orthogonal and parallel to the merger axis in the left and right panel, respectively. For the rest of this work, we adopt as the time shortly before the merger, visualized in the first row of Fig. 1. Almost directly after impact, two large shock shells emerge from the center, which together form an almost full circle with a diameter of after (third panel in second row of Fig. 1). While surpassing the diameter of ORC Physalis, these dimensions are generally agreeing well with the sizes of other observed ORCs, since Physalis is actually on the lower size end (c.f. Table 3 by Koribalski et al. 2024b). The clear boundary where the X-ray brightness is dropping off is coinciding with the position of the shock front, traced by the cosmic ray electron population in the simulation. The total magnitude of the X-ray flux is generally in agreement with observations of halos in this mass range (Bulbul et al. 2024). Meanwhile the merger shock is further expanding in radial direction, reaching a diameter of about after the merger event, hence far surpassing the virial radius of the system equal to .

### v2

@@ Abstract
Context. A new class of distinct radio objects, commonly referred to as odd radio circles (ORCs), has been recently discovered. The origin of these features remains unclear because their peculiar properties challenge our current understanding of astrophysical sources for diffuse radio emission.
@@
Methods. By modelling the assembly of a massive galaxy group with a final virial mass of in a magnetohydrodynamic zoom-in simulation with on-the-fly cosmic ray treatment, we derive the X-ray and radio properties of the system self-consistently and compare them to observations.
Results. We show that the X-ray properties of the simulated system agree with characteristics of observed galaxy groups in the relevant mass range, legitimating the comparison between the radio properties of the simulated halo and those of observed ORCs. A major merger between two galaxies in the simulation triggers a series of strong shocks in the circumgalactic medium, which in unison form a ring if the line of sight is perpendicular to the merger axis. The shock is rapidly expands radially and quickly reaches the virial radius of the halo. This formation channel thus readily explains the morphology and large extent of ORCs. However, the inferred radio luminosity of these features is lower than that of observed counterparts, while the degree of polarisation seems systematically over-predicted by the simulation.
Conclusions. Fossil cosmic ray populations from active galactic nuclei and stellar feedback might be necessary to explain the full extent of the radio properties of ORCs, since diffusive shock acceleration was the only source term for non-thermal electrons considered in this work.
@@
The current advent of next-generation radio telescopes delivers a number of interesting surprises to the astrophysics community. Their unprecedented depth and frequency coverage continue to reveal stunning details in various extraterrestrial settings—from galactic sources such as supernova remnants (e.g. Weiler et al. 1986; Dokara et al. 2021; Ball et al. 2023; Do et al. 2024) to the plentiful extragalactic zoo of objects such as star-forming disc galaxies (e.g. Martinsson et al. 2016; Krause et al. 2018; Stein et al. 2023; Heesen et al. 2024), galaxies hosting active galactic nuclei (AGN, e.g. McKean et al. 2021; Venturi et al. 2022; Koribalski et al. 2024a), as well as haloes and radio relics of galaxy clusters (e.g. van Weeren et al. 2010; Rajpurohit et al. 2020; Brüggen et al. 2021; Venturi et al. 2022; Cassano et al. 2023; Koribalski et al. 2024c). However, it seems that yet another type of distinct source needs to be added to the list of possible radio objects. As their name suggests, odd radio circles (ORCs) are peculiar radio features with edge-brightened ring- or shell-like morphologies (Norris et al. 2021b; Koribalski et al. 2021, 2024b). The spectral indices measured in these features, that can lie between and (assuming that the emission spectrum follows ), suggest that these regions are filled with shock-accelerated cosmic rays (CRs) emitting via synchrotron at radio frequencies. In agreement with other current studies on this topic, we only consider objects at high galactic latitudes that exhibit an identified galaxy close to their geometric centre, in order to separate potentially different phenomena (Koribalski et al. 2024b; Norris et al. 2025). Bona fide ORCs thus have a central early-type galaxy, strongly suggesting they are of extragalactic origin (Koribalski et al. 2021; Norris et al. 2022, 2025). Since ORCs lack a counterpart in any other wavelength regime, distance and size measurements of ORCs rely on the redshifts of these putative host galaxies. Judging by optical measurements of those objects, ORCs lie at distances of and often have diameters of several hundred kiloparsecs, while the stellar masses of the central galaxies lie around . This means that ORCs have enormous sizes compared to their central host, as their extent can even exceed the virial radius of the host system. While the majority of ORCs has been observed significantly above the Milky Way disc, we note that a few objects with similar morphologies have also been identified at low galactic latitudes (e.g. Koribalski et al. 2024c). But as mentioned above, they do not display a prominent central galaxy and differ in various other properties, indicating that these features are likely galactic supernova remnants and thus are formed in an entirely different setting (Filipović et al. 2022; Omar 2022a; Sarbadhicary et al. 2023).
@@
Since the nature of ORCs remains under debate, careful control of the sample of considered objects is needed to ensure an appropriate comparison. The following set of conditions seems to work best as a common denominator: the ORC must a) display a ring or shell-like morphology, b) host at least one elliptical galaxy in its centre, and c) have a large radio ring size comparable to the virial radius of the central galaxy (group). While the neighbouring ORCs 2 and 3 from the original discovery paper by Norris et al. (2021b) — who first reported such radio morphologies and have introduced the term ORC — likely have an AGN-related nature (e.g. Norris et al. 2021a, 2025; Shabala et al. 2024), the origin of a growing list of different ORCs remains unclear. It is also important to note that the observed morphologies can depend on the resolution and frequency of the observation facility. While an object may look diffuse in one image, the characteristic ring emission for ORCs can reveal itself in surveys with higher resolution and sensitivity.
The following observed objects are falling under the declared criteria: ORCs 1 and 4 (Norris et al. 2021b), ORC 5 (Koribalski et al. 2021), ORC 6 (Dolag et al. 2023; Koribalski et al. 2025), ORC J2223–4834 (Gupta et al. 2022), SAURON (Lochner et al. 2023), Physalis (Koribalski et al. 2024b), Cloverleaf (Bulbul et al. 2024), ORC J1027–4422 (Koribalski et al. 2024c), ORC J0219–0505 (Norris et al. 2025), ORC J0356–4216 (Taziaux et al. 2025) and RAD J131346.9+500320 (Hota et al. 2025) The benchmarks stated above for ORCs are rather restrictive and thus some sources published as ORCs are not included in our list. Nevertheless it is difficult to draw a definite line for candidates, as one cannot be sure of the exact set of properties they need to display111e.g. the diameters of Physalis (145 kpc) and ORC J0219–0505 (114 kpc) are relatively small compared to the other candidates (a few hundred kiloparsecs). Still the extent of their radio features is large compared to galaxy scales, and since all other properties fit very well with the rest of observed ORCs, we include them in our list.. We therefore build a quite conservative list, attempting to group only those objects likely to have a similar formation path. Nevertheless we note the several ongoing searches for ORC-like systems, as the study by Gupta et al. (2025).
@@
A variety of extragalactic formation scenarios for ORCs have been proposed. Several studies have addressed the scenario of a termination shock due to galactic outflows during starburst events (Norris et al. 2021b, 2022; Coil et al. 2024). Another possibility is that a black hole drives these features via binary mergers of supermassive black holes (Koribalski et al. 2021; Norris et al. 2022), tidal disruption of stars by black holes (Omar 2022b), or AGN jets, where the ejected radio lobes can appear circular under certain projections and physical conditions (Nolting et al. 2023; Fujita et al. 2024; Lin & Yang 2024). In this regard, Shabala et al. (2024) have pointed out that the radio power of a mere AGN jet would decline too quickly to explain the properties of ORCs and hence concluded that an additional shock running through pre-existing AGN bubbles would be necessary in this scenario. Finally, there is the shock-driven formation channel first proposed by Dolag et al. (2023) and Koribalski et al. (2024b). Based on cosmological simulations, the suggested concept posits that merger-accelerated shocks – i.e. merger shocks from coalescence events in massive galaxies and groups encountering outer accretion shocks – can produce features that closely resemble ORCs. Such processes have been studied for a while now in relation to galaxy clusters, where merger shocks are most likely the reason for radio relics (e.g. Vazza et al. 2009; Ha et al. 2018; Zhang et al. 2019). But in the case of clusters, these features always lie well within the boundaries of the respective halo. Hence, ORCs could be unprecedented evidence for merger shocks in galaxy and group haloes, where their extent pushes beyond the virial radius. However, the simulations by Dolag et al. (2023) predicted radio luminosities that were too low compared to observations, concluding that the relatively low simulated halo mass of could be a major issue, since the total energy budget is underestimated. This problem was also pointed out by Yamasaki et al. (2024), who analytically explored the feasibility of pure accretion shocks as the source for ORCs and concluded that more complex simulations with a larger halo mass of at least are required matching the estimated halo masses of ORC host galaxies (Norris et al. 2021a; Koribalski et al. 2024b).
@@
In the present work, we address the need for such a study. We present a first-of-its-kind zoom-in simulation of a galaxy merger in high resolution, where the final group mass corresponds to observed haloes in ORCs. We performed a non-radiative magnetohydrodynamic simulation with an on-the-fly treatment of spectral CR electrons. First, we summarise the numerical details of the simulation by stating the important features of the code and initial condition in Sect. 2. In Sect. 3, we give a detailed presentation of the simulation results. We outline the time evolution of the merger and then describe the halo’s X-ray and radio properties, including inferred mock observation images. Particular focus is placed on the polarisation of the simulated ORC. Similarities and differences with observations are discussed in Sect. 4, before we summarise our results in Sect. 5.
@@ 2 Simulation code and setup
The simulation was carried out with the cosmological code OpenGadget3, which utilises a smoothed particle magnetohydrodynamics (SPMHD) scheme to model gas dynamics and a Tree-PM solver for gravity (release paper by Dolag et al. in prep.). Its infrastructure is based on its predecessor Gadget2 (Springel 2005) but employs major improvements in the gravity solver (Ragagnin et al. 2016) and SPH implementation (Beck et al. 2016b), where also a revised time step criterion ensures the stability of the code around strong shocks in a simulation (Pakmor et al. 2012). We used the Wendland C4 function as the kernel to reconstruct the local hydrodynamic properties of the simulated fluid based on the position and smoothing length of the simulated particles (Dehnen & Aly 2012; Donnert et al. 2013). The code includes both a physically motivated conductivity (Jubelgas et al. 2004; Dolag et al. 2004; Petkova & Springel 2009), as well as an artificial conduction and time-dependent viscosity scheme (Price 2012; Beck et al. 2016b), which provides numerical convergence when compared to other hydrodynamic discretisation methods.
@@
The code resolves the magnetohydrodynamic equations in Lagrangian form (Dolag et al. 1999; Dolag & Stasyszyn 2009) in the non-ideal regime, i.e. with non-vanishing resistivity (Bonafede et al. 2011). Non-zero contributions from the magnetic field , arising from numerical truncation errors, are suppressed by the hyperbolic divergence cleaning method described by Tricco et al. (2016), whose implementation in OpenGadget3 is described by Steinwandel and Price (in prep.). We ran the code with an on-the-fly shock finder (Beck et al. 2016a), that returns the shock’s hydrodynamic quantities, as well as the Alfvénic Mach number and shock obliquity (Böss et al. 2023b). Finally, we includes CR electrons with the module CRESCENDO, a Fokker-Planck solver for spectral CRs (Böss et al. 2023b). The sub-grid model computes and evolves the distribution function of CRs for each resolution element (i.e. SPH particle) at each time step in the simulation. The source for CRs is diffusive shock acceleration (DSA), which heavily relies on the output of the built-in shock finder and uses shock-obliquity-dependent acceleration efficiencies (Pais et al. 2018; Kang 2024). These injected populations evolve further over time through adiabatic losses as well as radiative losses, the latter being inverse Compton scattering and synchrotron emission for CR electrons (for the details on the computation of the synchrotron emissivity, see Appendix 2 by Böss et al. 2023a). The present simulation was performed in a non-radiative regime without star formation, resolving only the dark matter and gas evolution.
@@
The initial condition for the simulation was drawn from the COMPASS222http://www.magneticum.org/complements.html#Compass set. It is a suite of zoom-in setups for a diverse selection of haloes – from massive clusters down to Milky Way-like galaxies – selected from a dark matter-only cosmological simulation with a box length of (e.g. Bonafede et al. 2011). We chose the dfrogin region, which produces a galaxy group halo with a final virial mass of (Schlachtberger 2014). The mass resolution for the gas and dark matter particles, as well as the corresponding smoothing lengths in the simulation are listed in Table 1. The simulation employs the standard WMAP-7 CDM cosmological parameters (Komatsu et al. 2011).
@@ 3 Results
Until , the halo undergoes an assembly process that includes various mergers with different galaxies. Shortly after , the main halo experiences a particularly violent event, with a mass ratio of . It is this major merger that triggers the shock and CR injection with subsequent radio emission that we identify as an ORC in this paper. In the following sections, we analyse the time evolution and properties of the galaxy group around this merger event. Where required, we adopt the distance to ORC Physalis while creating the mock maps and compare simulated properties to that observed counterpart. We acknowledge that it is not necessarily a representative case for the overall population of ORCs. However, given that it is found in the smallest observational volume with the least stringent observational constraints imposed by survey sensitivity, angular resolution and searching techniques, we consider it as a prototype for a large fraction of the ORC population333Physalis is the closest observed ORC so far and it would be highly unlikely to find a special case of an already rare phenomenon in such a small observational volume. This motivates the hypothesis that it represents a typical case, since ORCs with comparably small sizes and radio luminosities would not be detected at the much larger distances where other ORCs have been found.. Therefore, Physalis offers the best opportunities for a comparison between observational data and simulations at the moment.
@@ 3.1 Gas and shock morphology
Fig. 1 shows the evolution of the system around the time of the major merger event. The rows represent three different output times separated by . The left and right halves of the figure display the X-ray surface brightness in the band and the CR electron energy density, respectively. The X-ray flux was estimated based on contributions from thermal and line emission (assuming ), using the software Smac (Dolag et al. 2005a). We assumed a distance of (corresponding to ), as for the observed ORC Physalis (Koribalski et al. 2024b). We show two projections in each case, displaying the system orthogonal and parallel to the merger axis in the left and right panels, respectively. For the rest of this work, we adopt as the time shortly before the merger, visualised in the first row of Fig. 1. Almost directly after impact, two large shock shells emerge from the centre, which together form an almost full circle with a diameter of after (third panel in second row of Fig. 1). While surpassing the diameter of ORC Physalis, these dimensions generally agree well with the sizes of other observed ORCs, since Physalis is actually on the lower-size end (cf. Table 3 by Koribalski et al. 2024b). The clear boundary, where the X-ray brightness drops, coincides with the position of the shock front, traced by the CR electron population in the simulation. The total magnitude of the X-ray flux generally agrees with observations of haloes in this mass range (Bulbul et al. 2024). Meanwhile, the merger shock further expands in radial direction, reaching a diameter of about after the merger event, hence far surpassing the virial radius of the system equal to .